Shift report 20030223 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Feb 23 2003 - 16:47:23 EST

  • Next message: Apache: "Shift report 20030223 16:00-24:00"
    ------------------------------------------------------------
    Date: Sun Feb 23 16:47:22 2003, Shift: 08:00-16:00
    Supervisor: JHL, with: FR
    Log book#14 pages 33 - 33
    
    Shift Summary:
    ------------------------------------------------------------
    1. ~3 Hrs Beam with high background.
       TOFW/ZDC rate is 20-30 at MRS 90deg with ZDC~5-7K, 
       but large background in TPM1. 
       Information in TPM1 is unreconstructable.(See some
       "typical"  hits in TPM1 in the e-log)
       FS tracking seems okay.
    
    2. ~70K FS trigger at 12 degree (1/4B)
       (~150K MRS Trigger at 90 degree with hugh background 
        in TPM1)
    
    3. No detector problems except some old problems:
       -H2 slat7
       -T3 1Ay-u
      
     
       
        
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun Feb 23 2003 - 16:48:21 EST